Cleaning Status: Completed.
Findings: No problem found.
Cleaning Status: Completed.
Findings: No problem found.
106,117,118 এবং 223 এই DTW গুলো তে আগে পাওয়ার চলে গেলে ফল্ট ধরে রাখত যার কারণে অপরেশন টিম থেকে লোক পাঠিয়ে এটা রিসেট…
Problem: OGR Booster pump E running time Moter current high (682 Amp) showing than pump stop that time frequency is 0.3 Hz (Abnormal).Problem…
Due To Strom the optical fiber cable fall down and damage near DTW-121 to 122 now everything is normal
SL/NO PROBLEMS TAG LOCATION REMARK 1 When the pump runs, the pump Motor + MOV runs together. Which does not follow the normal…
NEV voltage [Report] SL No. Location Voltage Comments 01. DTW-101 37mV 02. DTW-102 170mV 03. DTW-103 231mV 04. DTW-104 1.1V 05. DTW-105 42mV…
One Comment